OpenStudy (anonymous):

How do you rewrite 2y^2+8y-2x+6=0 in standard form?

OpenStudy (is3535):

(y+2)^2=x-3

OpenStudy (anonymous):

How'd you get it?

OpenStudy (anonymous):

@is3535

OpenStudy (dan815):

Have you looked up what standard form is

OpenStudy (dan815):

2y^2+8y-2x+6=0 2x=2y^2+8y+6 x=y^2+4y+3